OpenStudy (anonymous):

Your cell phone company charges $50 per month for basic service, plus an additional $0.10 per text message. Let m represent the number of text messages and C represent the total monthly bill, write an equation to find the total monthly bill from your cell phone company

OpenStudy (anonymous):

total = $50 + m*(cost of 1 message) so the equation is : C = 50 + 0.1m => 10C= 500 + m so C = 500+m over 10 (500+m /10) [this gives total cost in a month ] Hope I could help. :)